Caffeine is a stimulant found in coffee and has chemical formula C8H10N4O2 . A 100g sample of coffee powder contains 980 mg caffeine. Determine %nitrogen in 42.0 g of coffee powder

1
Expert's answer
2022-06-06T23:32:06-0400

Mr (C8H10N4O2) = 194 g/mol

1 mol (194 g) of caffeine contains 4 atoms (56 g) N.


1.

100 g of coffee contains 980 mg of caffeine. Find how much caffeine is in 42 g of coffee:

100 g coffee — 980 mg caffeine

42 g coffee — X

X = 42 * 980 / 100 = 411,6 g caffeine


2.

194 g of caffeine contains 56 g of N. We find how much N is in 411,6 mg of caffeine:

194 g caffeine — 56 g N

411,6 mg caffeine — X

X = 411,6 * 56 / 194 = 118,8 mg =0,1188 g N


3.

Find the percentage of nitrogen in 42 g of coffee. To do this, we divide the mass of nitrogen by the mass of coffee and multiply by 100:

0,1188 / 42 * 100% = 0,283 %

AMSWER: 0,238%